Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Question from GmatPrep Test

Expert replies
by onlyGmat2008 » Tue May 27, 2008 9:07 am
Hello,

I am not able to solve the problem (attached).

Could you please help me with this?

Thanks.
You do not have the required permissions to view the files attached to this post.
Join the discussion
Source: — Problem Solving |

by VP_Tatiana » Tue May 27, 2008 8:16 pm
Ok, so we have 15 slate rocks out of 45 rocks total.

The first time we pick a rock, we have a 15/45 = 1/3 chance of getting a slate rock.

Since we are not replacing that rock, we now have 14 slate rocks and 44 rocks total, for a 14/44 = 7/22 chance of getting a slate rock.

In order to get two slate rocks, we need to be in the lucky 1/3 that get a slate rock the first time, and then in the lucky 7/22 that get a slate rock the second time. To calculate the probability of being lucky both times, we need to multiply these outcomes together.

1/3 * 7/22 = 7/66

Hope that helps,

Tatiana
Tatiana Becker | GMAT Instructor | Veritas Prep
Join the discussion

by onlyGmat2008 » Tue May 27, 2008 8:50 pm
Thank You so much.
Join the discussion

by manu217 » Wed Jun 11, 2008 3:12 am
Thank you for your reply. However could you please clarify why did you multiply 1/3 and 7/22 and not add the probabilities?
Thanks
Join the discussion